Suppose that a laboratory test found the combined antibiotic effect of x2ri73+++sX2i7Uscr0iIUAAAAASUVORK5CYII= milligrams of medicine A and ymLxA6DwostvpbZrAAAAAElFTkSuQmCC milligrams of medicine B to be given by the function: LJjlyGBQt2YAAAAASUVORK5CYII=

  1. Find the amounts of the two medicines that maximize the antibiotic effect. You may assume concavity. (5 pts)
Answer #1

Answer:

Given Combined antibiotic effect of x milligrams (05x555) of medicine A and y milligrams (OS S60) of medicine B is represente

Differentiating the function f(x,y) partially with respect to x and y gives DE 1(x,y) =-4x+y+110; D = S(x,y) = x-2y+60 Solvin

The second order partial derivatives of the function P(x, y) are given by

  1585251085629_image.png

Since the second order partial derivatives are all constant, applying D-test to the above values leads to

D= fx19-[1.] = 8-1

Since D>0 and fa<0, by D-test the function f(x,y) has maximum value for x = 40 and y = 50